OpenStudy (anonymous):

I award metals! I need help please! I thought I knew the anser but I dont think I do now

OpenStudy (anonymous):

Jordan is a manager of a car dealership. He has 3 professional car washers to clean the entire lot of cars, Jennifer, Arianna, and Matthew. Jennifer can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Matthew is new to the car dealership, so no one knows how long it will take him. Hayden assigns all of them to wash the cars together. Explain to Jordan how this task can tell her how long it would take Matthew to complete the task if he worked by himself. Use complete sentences.

OpenStudy (anonymous):

Well first you add 1/11 and 1/14 by finding the common denominator and correcting the fraction. When you do that, you should get 14/ 154 + 14/154 = 28/154 then you cross multiply 28/154 and 1/x. Is that right?

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@Diyadiya

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

\[\frac{1}{14}+\frac{1}{11}+\frac{1}{m}=\frac{1}{t} \]Solve the above equation for m after replacing t with the measured total wash time required with all three washers working together.